# Model-free policy gradient for discrete-time mean-field control

### v1

@@ 5.1 Proofs of Results of Section 2
$\mathrm{MD}_{\varepsilon}(\theta)$ $\coloneqq\mathbb{E}\left[\left(\sum_{t=0}^{T-1}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(X_{T}^{\theta},l^{\theta}_{T}+\varepsilon\Lambda_{T})\right)\cdot\left(\sum_{s=0}^{T}\varepsilon^{-1}\Lambda_{s}\nabla_{\theta}l_{s}^{\theta}\right)\right],$
$\mathrm{MFD}_{\varepsilon}(\theta)$ $\coloneqq\mathbb{E}\Bigg[\left(\sum_{t=0}^{T-1}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(X_{T}^{\theta},l^{\theta}_{T}+\varepsilon\Lambda_{T})\right)\cdot\left(\sum_{s=0}^{T}\Lambda_{s}\nabla_{\theta}l_{s}^{\theta}\right)$
$\cdot\left(\sum_{\tau=0}^{T-1}\left(\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})+\nabla_{l}\log\tilde{p}(\theta,\alpha_{\tau}^{\theta},\tau,X_{\tau}^{\theta},l_{\tau}^{\theta})\right)\cdot\Lambda_{\tau}\right)\Bigg].$
The off-diagonal terms in $\mathrm{MD}_{\varepsilon}(\theta)$ cancel by independence of $(\Lambda_{t})_{t=0}^{T-1}$, hence
@@
$=\mathbb{E}\left[\varepsilon^{-1}(\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})-\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}))\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}\right]$
$=\mathbb{E}\left[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})\cdot\Lambda_{t}\left(\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}\right)\right]+o(1)$
$=\mathbb{E}[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})]\mathbb{E}[\Lambda_{t}^{T}\Lambda_{t}]\nabla_{\theta}l_{t}^{\theta}+o(1)$
$=\mathbb{E}[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})]\nabla_{\theta}l_{t}^{\theta}+o(1).$
@@
$\mathbf{r}_{\varepsilon}(\theta)=\sum_{t=0}^{T-1}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(X_{T}^{\theta},l^{\theta}_{T}+\varepsilon\Lambda_{T}).$
Note that $\mathbb{E}[\mathbf{r}_{\varepsilon}(\theta)\Lambda_{s}^{T}\Lambda_{\tau}]=0$ whenever $s\neq\tau$ and that $\mathbf{r}_{\varepsilon}(\theta)$ is a random variable converging almost surely to $\mathbf{r}(\theta)\coloneqq\mathbf{r}_{0}(\theta)$, hence
$\mathrm{MFD}_{\varepsilon}(\theta)$ $=\mathbb{E}\left[\mathbf{r}_{\varepsilon}\sum_{s=0}^{T-1}\left(\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})+\nabla_{l}\log\tilde{p}(\theta,\alpha_{\tau}^{\theta},\tau,X_{\tau}^{\theta},l_{\tau}^{\theta})\right)\cdot\Lambda_{s}\left(\Lambda_{s}\nabla_{\theta}l_{t}^{\theta}\right)\right]$
$=\mathbb{E}\left[\mathbf{r}\sum_{s=0}^{T-1}\left(\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})\tilde{p}(\theta,\alpha_{\tau}^{\theta},\tau,X_{\tau}^{\theta},l_{\tau}^{\theta})\right)\cdot\Lambda_{s}\left(\Lambda_{s}\nabla_{\theta}l_{t}^{\theta}\right)\right]+o(1)$
$=\mathbb{E}\left[\mathbf{r}\sum_{s=0}^{T-1}\left(\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})\tilde{p}(\theta,\alpha_{\tau}^{\theta},\tau,X_{\tau}^{\theta},l_{\tau}^{\theta})\right)\nabla_{\theta}l_{t}^{\theta}\right]+o(1),$
@@ 5.2 Proofs of Results of Section 3
Let $t\in\{0,\ldots,T\}$ and $i\in\{1,\ldots,d\}$. Denote $\mu_{t}^{\theta,i}\coloneqq\mathbb{P}(X_{t}^{\theta}=x^{(i)})$. We set the following notation, for all $x^{\prime},x\in\mathcal{X},a\in\mathcal{A},l\in{\mathbb{R}}^{d},\theta\in\Theta$,
$\varphi^{(i)}(\theta,x^{\prime},a,t,x,l)\coloneqq\mathbf{1}_{x^{\prime}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,a,t,x,l),\quad\psi(\theta,x^{\prime},t,x,a,l)\coloneqq\tilde{P}(x^{\prime}\mid x,a,l)\tilde{p}(\theta,a,t,x,l).$
By Proposition 2.1 with reward function $x\mapsto 1_{x=x^{(i)}}$ at time $t$ and 0 elsewhere,
@@
$=\Bigg\|\int\mathbf{1}_{x_{t}=x^{(i)}}\sum_{s=0}^{t-1}\frac{\nabla_{l}\psi(\theta,x_{s+1},s,x_{s},a_{s},l_{s}^{\theta})}{\psi(\theta,x_{s+1},s,x_{s},a_{s},l_{s}^{\theta})}\nabla_{\theta}l_{s}^{\theta}$
$\quad\quad\cdot\left(\prod_{\tau=0}^{t-1}\psi(\theta,x_{\tau+1},\tau,x_{\tau},a_{\tau},l_{\tau}^{\theta})\nu_{\mathcal{X}}({\mathrm{d}}x_{\tau+1})\nu_{\mathcal{A}}({\mathrm{d}}a_{\tau})\right)\mu_{0}({\mathrm{d}}x_{0})\Bigg\|_{\infty}$
$\leq\int\mathbf{1}_{x_{t}=x^{(i)}}\sum_{s=0}^{t-1}B_{1}\left\|\nabla_{\theta}l_{s}^{\theta}\right\|$
$\quad\quad\cdot\prod_{\begin{subarray}{c}\tau=0\\ \tau\neq s\end{subarray}}^{t-1}\psi(\theta,x_{\tau+1},\tau,x_{\tau},a_{\tau},l_{\tau}^{\theta})\nu_{\mathcal{X}}({\mathrm{d}}x_{\tau+1})\nu_{\mathcal{A}}({\mathrm{d}}a_{\tau})\nu_{\mathcal{X}}({\mathrm{d}}x_{s+1})\nu_{\mathcal{A}}({\mathrm{d}}a_{s})\mu_{0}({\mathrm{d}}x_{0})$
$\leq\left\|\nabla_{l}\psi\right\|_{\infty}|\mathcal{A}|\sum_{s=0}^{t-1}d^{\mathbf{1}_{s<t-1}}\left\|\nabla l_{s}^{\theta}\right\|.$
@@
$\mathbf{s}^{\theta}\coloneqq\sup_{0\leq t\leq T}\sum_{i=1}^{d}\mathbb{P}(X_{t}^{\theta}=x^{(i)})^{-1}.$
@@ 5.2.1 Proofs on Estimation Biases
Let $\theta\in\Theta$ and fix it for the rest of the proof. For the whole proof, we denote by $X_{t}^{\theta},Y_{t}^{\theta,\varepsilon},\alpha_{t}^{\theta},\alpha_{t}^{\theta,\varepsilon}$ processes distributed according to (2.3) using perturbations $\Lambda_{t},t=0,\ldots,T-1$, independent of the processes and perturbations used in the estimators $\widehat{\nabla_{\theta}l_{t}^{\theta}}$. The bound is trivial for $t=0$. We first assume $B_{1}>0$. Let $t\geq 1$. For $i=1,\ldots,d$, let $\mu^{(i)}_{t}=\mathbb{P}(X_{t}^{\theta}=x^{(i)})$. We set the following notations for the rest of the proof:
$E_{t}(\theta,\varepsilon)\coloneqq\left\|\mathbb{E}\left[\widehat{\nabla_{\theta}l_{t}^{\theta}}\right]-\nabla_{\theta}l_{t}^{\theta}\right\|,\>b_{t}^{(i)}(\theta)\coloneqq\mathbb{E}\left[\widehat{\nabla_{\theta}\mu_{t}^{(i)}}\right]-\nabla_{\theta}\mu^{(i)}_{t}$. (5.5)
@@
$+\left(\nabla_{l}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})+\nabla_{l}\log\tilde{P}(X_{s+1}^{\theta}\mid X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\right)\nabla_{\theta}l_{s}^{\theta}\Big)\Bigg].$
@@
$b_{t,1}^{(i)}(\theta)$ $\coloneqq\mathbb{E}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\sum_{s=0}^{t-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\varepsilon\Lambda_{s})\right]$
@@
$b_{t,2}^{(i)}(\theta)$ $\coloneqq\mathbb{E}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\sum_{s=0}^{t-1}\varepsilon^{-1}\Lambda_{s}\widehat{\nabla_{\theta}l_{s}^{\theta}}\right$]
@@
$\Delta_{1}^{(i)}(\theta)$ $\coloneqq\mathbb{E}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\sum_{s=0}^{t-1}\left(\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\varepsilon\Lambda_{s})-\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta})\right)\right],$
$\Delta_{2}^{(i)}(\theta)$ $\coloneqq\mathbb{E}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\sum_{s=0}^{t-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta})-\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\sum_{s=0}^{t-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta}).\right]$
Since $\nabla_{\theta}p$ is Lipschitz in the measure argument, we have
@@
$R_{t}^{\theta,\varepsilon}$ $\coloneqq\prod_{s=0}^{t-1}\frac{\tilde{P}(X^{\theta}_{s+1}\mid X^{\theta}_{s},\alpha_{s}^{\theta},l_{s}^{\theta}+\varepsilon\Lambda_{s})\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta}+\varepsilon\Lambda_{s})}{\tilde{P}(X^{\theta}_{s+1}\mid X^{\theta}_{s},\alpha_{s}^{\theta},l_{s}^{\theta})\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})}$ (5.7)
@@
$+\frac{\varepsilon^{2}}{2}\frac{\sum_{j,k=1}^{d}\xi^{\theta}_{s_{r},j,k}\Lambda_{s_{r}}^{(j)}\Lambda_{s_{r}}^{(k)}}{\psi(\theta,X_{s_{r}+1}^{\theta},X_{s_{r}}^{\theta},\alpha_{s_{r}}^{\theta},l_{s_{r}}^{\theta})}\Bigg).$
@@
$U^{\theta}_{\tau}$ $=\frac{\sum_{j,k=1}^{d}\xi^{\theta}_{\tau,j,k}\Lambda_{\tau}^{(j)}\Lambda_{\tau}^{(k)}}{\psi(\theta,X_{\tau+1}^{\theta},X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})},$
@@
$\left\|\|\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\Lambda_{s}\sum_{\tau=0}^{t-1}U_{\tau}^{\theta}\right]\right\|_{\infty}=\left\|\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\Lambda_{s}U_{s}^{\theta}\right]\right\|_{\infty}\leq B_{2}d^{3}|\mathcal{A}|\mathbb{E}[|Z|^{3}],$
@@
$=\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\Lambda_{s}\left(\varepsilon\nabla_{l}\log\psi(\theta,X^{\theta}_{s+1},s,X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}\prod_{r=1,s_{r}\neq s}^{k}\frac{\varepsilon^{2}}{2}U_{s_{r}}^{\theta}+\prod_{r=1}^{k}\frac{\varepsilon^{2}}{2}U^{\theta}_{s_{r}}\right)\right],$
@@
$\left\|\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\Lambda_{s}\nabla_{l}\log\psi(\theta,X^{\theta}_{s+1},s,X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}\prod_{r=1,s_{r}\neq s}^{k}\frac{\varepsilon^{2}}{2}U^{\theta}_{s_{r}}\right]\right\|_{\infty}$
$\leq\frac{\varepsilon^{2k-1}}{2^{k-1}}B_{1}B_{2}^{k-1}d^{2(k-1)}d^{k}|\mathcal{A}|^{k}\mathbb{E}[|Z|^{2k}],$
@@
Denote by $\sigma_{k}\coloneqq\mathbb{E}|Z|^{k}$. Summing over $k$ and all possible choices of $s_{1},\ldots,s_{k}$ including $s$, we get the following bound
@@
$+\sum_{k=2}^{t-1}\binom{t-1}{k-1}\left(\frac{\varepsilon^{2k-1}}{2^{k-1}}B_{1}B_{2}^{k-1}d^{2(k-1)}d^{k}|\mathcal{A}|^{k}\sigma_{2k}+\frac{\varepsilon^{2k}}{2^{k}}B_{2}^{k}d^{3k}|\mathcal{A}|^{k}\sigma_{2k+1}\right),$

### v2

@@ 5.1 Proofs of Results of Section 2
$\mathrm{MD}_{\varepsilon}(\theta)$ $\coloneq\mathbb{E}\left[\left(\sum_{t=0}^{T-1}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(X_{T}^{\theta},l^{\theta}_{T}+\varepsilon\Lambda_{T})\right)\cdot\left(\sum_{s=0}^{T}\varepsilon^{-1}\Lambda_{s}\nabla_{\theta}l_{s}^{\theta}\right)\right],$
$\mathrm{MFD}_{\varepsilon}(\theta)$ $\coloneq\mathbb{E}\Bigg[\left(\sum_{t=0}^{T-1}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(X_{T}^{\theta},l^{\theta}_{T}+\varepsilon\Lambda_{T})\right)\cdot\left(\sum_{s=0}^{T}\Lambda_{s}\nabla_{\theta}l_{s}^{\theta}\right)$
$\quad\cdot\left(\sum_{\tau=0}^{T-1}\left(\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})+\nabla_{l}\log\tilde{p}(\theta,\alpha_{\tau}^{\theta},\tau,X_{\tau}^{\theta},l_{\tau}^{\theta})\right)\cdot\Lambda_{\tau}\right)\Bigg].$
The off-diagonal terms in $\mathrm{MD}_{\varepsilon}(\theta)$ cancel by independence of $(\Lambda_{t})_{t=0}^{T}$, hence
@@
$\qquad=\mathbb{E}\left[\varepsilon^{-1}(\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})-\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}))\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}\right]$
$\qquad=\mathbb{E}\left[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})\cdot\Lambda_{t}\left(\Lambda_{t}\nabla_{\theta}l_{t}^{\theta}\right)\right]+o(1)$
$\qquad=\mathbb{E}[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})]\mathbb{E}[\Lambda_{t}^{T}\Lambda_{t}]\nabla_{\theta}l_{t}^{\theta}+o(1)$
$\qquad=\mathbb{E}[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})]\nabla_{\theta}l_{t}^{\theta}+o(1).$
@@
$\tilde{G}^{\varepsilon,\theta}_{0}=\sum_{t=0}^{T-1}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})+\tilde{g}(X_{T}^{\theta},l^{\theta}_{T}+\varepsilon\Lambda_{T}).$
Note that $\mathbb{E}[\tilde{G}^{\varepsilon,\theta}(\theta)\Lambda_{s}^{T}\Lambda_{\tau}]=0$ whenever $s\neq\tau$ and that $\tilde{G}^{\varepsilon,\theta}(\theta)$ is a random variable converging almost surely to $G_{0}^{\theta}$, hence
$\mathrm{MFD}_{\varepsilon}(\theta)$ $=\mathbb{E}\left[\tilde{G}^{\varepsilon,\theta}_{0}\sum_{s=0}^{T-1}\left(\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})+\nabla_{l}\log\tilde{p}(\theta,\alpha_{\tau}^{\theta},\tau,X_{\tau}^{\theta},l_{\tau}^{\theta})\right)\cdot\Lambda_{s}\left(\Lambda_{s}\nabla_{\theta}l_{t}^{\theta}\right)\right]$
$=\mathbb{E}\left[G_{0}^{\theta}\sum_{s=0}^{T-1}\left(\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})\tilde{p}(\theta,\alpha_{\tau}^{\theta},\tau,X_{\tau}^{\theta},l_{\tau}^{\theta})\right)\cdot\Lambda_{s}\left(\Lambda_{s}\nabla_{\theta}l_{t}^{\theta}\right)\right]+o(1)$
$=\mathbb{E}\left[G_{0}^{\theta}\sum_{s=0}^{T-1}\left(\nabla_{l}\log\tilde{P}(X_{\tau+1}^{\theta}\mid X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})\tilde{p}(\theta,\alpha_{\tau}^{\theta},\tau,X_{\tau}^{\theta},l_{\tau}^{\theta})\right)\nabla_{\theta}l_{t}^{\theta}\right]+o(1),$
@@ 5.2 Proofs of Results of Section 3
Let $t\in\{0,\ldots,T\}$ and $i\in\{1,\ldots,d\}$. We set the following notation, for all $x^{\prime},x\in\mathcal{X},a\in\mathcal{A},l\in{\mathbb{R}}^{d},\theta\in\Theta,s\in\{0,\ldots,T-1\}$,
$\varphi^{(i)}(\theta,x^{\prime},a,t,x,l)\coloneq\mathbf{1}_{x^{\prime}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,a,t,x,l),\quad\psi(\theta,x^{\prime},t,x,a,l)\coloneq\tilde{P}(x^{\prime}\mid x,a,l)\tilde{p}(\theta,a,t,x,l).$
Let $\nabla_{\theta}\mu_{t}^{(i)}=\nabla_{\theta}\mathbb{P}(X_{t}^{\theta}=x^{(i)})$. By Proposition 2.1 with reward function $x\mapsto 1_{x=x^{(i)}}$ at time $t$ and 0 elsewhere,
@@
$\qquad=\Bigg\|\int\mathbf{1}_{x_{t}=x^{(i)}}\sum_{s=0}^{t-1}\frac{\nabla_{l}\psi(\theta,x_{s+1},s,x_{s},a_{s},l_{s}^{\theta})}{\psi(\theta,x_{s+1},s,x_{s},a_{s},l_{s}^{\theta})}\nabla_{\theta}l_{s}^{\theta}$
$\qquad\quad\cdot\left(\prod_{\tau=0}^{t-1}\psi(\theta,x_{\tau+1},\tau,x_{\tau},a_{\tau},l_{\tau}^{\theta})\nu_{\mathcal{X}}({\mathrm{d}}x_{\tau+1})\nu_{\mathcal{A}}({\mathrm{d}}a_{\tau})\right)\mu_{0}({\mathrm{d}}x_{0})\Bigg\|_{\infty}$
$\qquad\leq\int\mathbf{1}_{x_{t}=x^{(i)}}\sum_{s=0}^{t-1}B_{1}\left\|\nabla_{\theta}l_{s}^{\theta}\right\|$
$\qquad\quad\cdot\prod_{\begin{subarray}{c}\tau=0\\ \tau\neq s\end{subarray}}^{t-1}\psi(\theta,x_{\tau+1},\tau,x_{\tau},a_{\tau},l_{\tau}^{\theta})\nu_{\mathcal{X}}({\mathrm{d}}x_{\tau+1})\nu_{\mathcal{A}}({\mathrm{d}}a_{\tau})\nu_{\mathcal{X}}({\mathrm{d}}x_{s+1})\nu_{\mathcal{A}}({\mathrm{d}}a_{s})\mu_{0}({\mathrm{d}}x_{0})$
$\qquad\leq\left\|\nabla_{l}\psi\right\|_{\infty}|\mathcal{A}|\sum_{s=0}^{t-1}d^{\mathbf{1}_{s<t-1}}\left\|\nabla l_{s}^{\theta}\right\|.$
@@
$\mathbf{s}^{\theta}\coloneq\sup_{0\leq t\leq T}\sum_{i=1}^{d}\mathbb{P}(X_{t}^{\theta}=x^{(i)})^{-1}.$
@@ 5.2.1 Proofs on Estimation Biases
Let $\theta\in\Theta$ and fix it for the rest of the proof. For the whole proof, we denote by $X_{t}^{\theta},Y_{t}^{\theta,\varepsilon},\alpha_{t}^{\theta},\alpha_{t}^{\theta,\varepsilon}$ processes distributed according to (2.3) using perturbations $\Lambda_{t},t=0,\ldots,T$, independent of the processes and perturbations used in the estimators $\widehat{\nabla_{\theta}l_{t}^{\theta}}$. The bound is trivial for $t=0$. We first assume $B_{1}>0$. Let $t\geq 1$. For $i=1,\ldots,d$, let $\mu^{(i)}_{t}=\mathbb{P}(X_{t}^{\theta}=x^{(i)})$. We set the following notations for the rest of the proof:
$E_{t}(\theta,\varepsilon)\coloneq\left\|\mathbb{E}\left[\widehat{\nabla_{\theta}l_{t}^{\theta}}\right]-\nabla_{\theta}l_{t}^{\theta}\right\|,\>b_{t}^{(i)}(\theta)\coloneq\mathbb{E}\left[\widehat{\nabla_{\theta}\mu_{t}^{(i)}}\right]-\nabla_{\theta}\mu^{(i)}_{t}$. (5.5)
@@
$\quad+\left(\nabla_{l}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})+\nabla_{l}\log\tilde{P}(X_{s+1}^{\theta}\mid X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\right)\nabla_{\theta}l_{s}^{\theta}\Big)\Bigg].$
@@
$b_{t,1}^{(i)}(\theta)$ $\coloneq\mathbb{E}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\sum_{s=0}^{t-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\varepsilon\Lambda_{s})\right]$
@@
$b_{t,2}^{(i)}(\theta)$ $\coloneq\mathbb{E}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\sum_{s=0}^{t-1}\varepsilon^{-1}\Lambda_{s}\widehat{\nabla_{\theta}l_{s}^{\theta}}\right$]
@@
$\Delta_{1}^{(i)}(\theta)$ $\coloneq\mathbb{E}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\sum_{s=0}^{t-1}\left(\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta}+\varepsilon\Lambda_{s})-\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta})\right)\right],$
$\Delta_{2}^{(i)}(\theta)$ $\coloneq\mathbb{E}\left[\mathbf{1}_{Y_{t}^{\theta,\varepsilon}=x^{(i)}}\sum_{s=0}^{t-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta,\varepsilon},s,Y_{s}^{\theta,\varepsilon},l_{s}^{\theta})-\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\sum_{s=0}^{t-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta}).\right]$
Since $\nabla_{\theta}\log p$ is Lipschitz in the measure argument, we have
@@
$R_{t}^{\theta,\varepsilon}$ $\coloneq\prod_{s=0}^{t-1}\frac{\tilde{P}(X^{\theta}_{s+1}\mid X^{\theta}_{s},\alpha_{s}^{\theta},l_{s}^{\theta}+\varepsilon\Lambda_{s})\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta}+\varepsilon\Lambda_{s})}{\tilde{P}(X^{\theta}_{s+1}\mid X^{\theta}_{s},\alpha_{s}^{\theta},l_{s}^{\theta})\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})}$ (5.7)
@@
$\quad+\frac{\varepsilon^{2}}{2}\frac{\sum_{j,k=1}^{d}\xi^{\theta}_{s_{r},j,k}\Lambda_{s_{r}}^{(j)}\Lambda_{s_{r}}^{(k)}}{\psi(\theta,X_{s_{r}+1}^{\theta},s_{r},X_{s_{r}}^{\theta},\alpha_{s_{r}}^{\theta},l_{s_{r}}^{\theta})}\Bigg).$
@@
$U^{\theta}_{\tau}$ $=\frac{\sum_{j,k=1}^{d}\xi^{\theta}_{\tau,j,k}\Lambda_{\tau}^{(j)}\Lambda_{\tau}^{(k)}}{\psi(\theta,X_{\tau+1}^{\theta},\tau,X_{\tau}^{\theta},\alpha_{\tau}^{\theta},l_{\tau}^{\theta})},$
@@
$\left\|\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\Lambda_{s}\sum_{\tau=0}^{t-1}U_{\tau}^{\theta}\right]\right\|_{\infty}=\left\|\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\Lambda_{s}U_{s}^{\theta}\right]\right\|_{\infty}\leq B_{2}d^{3}|\mathcal{A}|\mathbb{E}[|Z|^{3}],$
@@
$\qquad=\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\Lambda_{s}\left(\varepsilon\nabla_{l}\log\psi(\theta,X^{\theta}_{s+1},s,X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}\prod_{r=1,s_{r}\neq s}^{k}\frac{\varepsilon^{2}}{2}U_{s_{r}}^{\theta}+\prod_{r=1}^{k}\frac{\varepsilon^{2}}{2}U^{\theta}_{s_{r}}\right)\right],$
@@
$\left\|\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\Lambda_{s}\varepsilon\nabla_{l}\log\psi(\theta,X^{\theta}_{s+1},s,X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}\prod_{r=1,s_{r}\neq s}^{k}\frac{\varepsilon^{2}}{2}U^{\theta}_{s_{r}}\right]\right\|_{\infty}$
$\qquad\leq\frac{\varepsilon^{2k-1}}{2^{k-1}}B_{1}B_{2}^{k-1}d^{2(k-1)}d^{k}|\mathcal{A}|^{k}\mathbb{E}[|Z|^{2k}],$
@@
Denote by $\sigma_{k}\coloneq\mathbb{E}|Z|^{k}$. Summing over $k$ and all possible choices of $s_{1},\ldots,s_{k}$ including $s$, we get the following bound
@@
$\quad+\sum_{k=2}^{t-1}\binom{t-1}{k-1}\left(\frac{\varepsilon^{2k-1}}{2^{k-1}}B_{1}B_{2}^{k-1}d^{2(k-1)}d^{k}|\mathcal{A}|^{k}\sigma_{2k}+\frac{\varepsilon^{2k}}{2^{k}}B_{2}^{k}d^{3k}|\mathcal{A}|^{k}\sigma_{2k+1}\right),$